40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 30 00:02:25 2013, Jenne, Update, LSC, LSC demod boards need some thought 
    Reply  Wed Oct 30 01:22:56 2013, Jenne, Update, LSC, REFL 165 demod phase adjusted 
       Reply  Wed Oct 30 01:44:13 2013, Jenne, Update, LSC, MICH and PRCL gains adjusted (Config file saved) 
          Reply  Wed Oct 30 03:36:51 2013, Jenne, Update, LSC, MICH and PRCL UGFs change with ALS enabled 
             Reply  Mon Nov 4 15:46:17 2013, Jenne, Update, LSC, Thoughts and Conclusions from last week's PRMI+2arms attempt 12x
                Reply  Tue Nov 5 19:39:27 2013, Jenne, Update, LSC, OpLev time series OpLevs_during_PRMI_2arms.pdf
    Reply  Wed Oct 30 03:33:17 2013, Rana, Update, LSC, LSC demod boards need some thought 0309.png
       Reply  Wed Oct 30 08:32:55 2013, Evan, HowTo, Computer Scripts / Programs, matplotlibrc 
Message ID: 9338     Entry time: Mon Nov 4 15:46:17 2013     In reply to: 9317     Reply to this: 9349
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Thoughts and Conclusions from last week's PRMI+2arms attempt 

5:31pm - This is still a work in progress, but I'm going to submit so that I save my writing so far. I think I'm done writing now.


First, a transcription of some of the notes that I took last Tuesday night, then a few looks at the data, and finally some thoughts on things to investigate.


MICH and PRCL Transfer Functions while arms brought in to resonance (both arms locked to ALS beatnotes):

This is summarized in elog 9317, which I made as we were finishing up Tuesday night.  Here's the full story though.  Note that I didn't save the data for these, I just took notes (and screenshots for the 1st TF).

POP22I was ~140 counts, POP110I was ~100 counts.

MICH gain = -2.0, PRCL gain = 0.070. 

First TF (used as reference for 2-10), PRMI locked on REFL165, Xarm transmission = 0.03, Yarm transmission = 0.05 (both arms off resonance).  MICH UGF~40Hz, PRCL UGF~80Hz.

MICH_40Hz.pngPRCL_80Hz.png

2: X=off-res (xarm not moved), Y=0.13, no change in TF

3: X=off-res (xarm not moved), Y=0.35, no change in TF

4: X=off-res (xarm not moved), Y=0.60, MICH high freq gain went up a little, otherwise no change (no change in either UGF)

5:  X=off-res (xarm not moved), Y=0.95, same as TF#4.

6: X=0.20, Y=1.10 (yarm not moved), same as TF#4

7:  X=0.40, Y=1.30 (yarm not moved), same as TF#4

8:  X=0.70, Y=1.55 (yarm not moved), same as TF#4

9: X=1.40, Y=2.20 (yarm not moved), same as TF#4

10: X=4.0, Y=4.0 (yarm not moved), PRCL UGF is 10Hz higher than TF#4, MICH UGF is 20Hz lower than TF#4.

11: (No TF taken), Xarm and Yarm transmission both around 20!  To get this, MICH FMs that were triggered, are no longer triggered to turn on.  Also, MICH gain was lowered to -0.15 and PRCL gain was increased to 0.1

12: (No TF taken), Xarm and Yarm transmissions both around 40!  The peaks could be higher, but we don't have the QPD ready yet.

After that, we started moving away from resonance, but we didn't take any more transfer functions.


OpLev spectra for different arm resonance values:

We were concerned that the ETMs and ITMs might be moving more, when the arms are resonating high power, due to some optical spring / radiation pressure effects, so I took spectra of oplevs at various arm transmissions.

I titled the first file "no lock", and unfortunately I don't remember what wasn't locked.  I think, however, that nothing at all was locked.  No PRMI, no arm ALS, no nothing.  Anyhow, here's the spectrum:

ALS_noLock.pdf

I have a measurement when the Yarm's transmission was 1, and the Xarm's transmission was 1.75.  This was a PRMI lock, with ALS holding the arms partially on resonance:

ALS_X1pt75Y1.pdf

Next up, I have a measurement when Yarm was 0.8, Xarm was 2.  Again, PRMI with the arms held by ALS:

ALS_X2Y0pt8.pdf

And finally, a measurement when Xarm was 5, Yarm was 4:

ALS_X5Y4.pdf

Just so we have a "real" reference, I have just now taken a set of oplev spectra, with the ITMs, ETMs and PRM restored, but I shut the PSL shutter, so there was no light flashing around pushing on things.  I noticed, when taking this data, that if the PSL shutter was open, so the PRFPMI is flashing (but LSC is off), the PRM oplev looks much like the original "no Lock" spectra, but when I closed the shutter, the oplev looks like the others.  So, perhaps when we're getting to really high powers, the PRM is getting pushed around a bit?

ALS_noLock_noLaser.pdf

Conclusions from OpLev Spectra:  At least up to these resonances (which is, admittedly, not that much), I do not see any difference in the oplev spectra at the different buildup power levels.  What I need to do is make sure to take oplev spectra next time we do the PRMI+2arms test when the arms are resonating a lot. 


Time series while bringing arms into resonance:

PRMI_2arms_29Oct2013_POPrin.png

I had wondered if, since the POP 22 and 110 values looked so shakey, we were increasing the PRCL RIN while we brought the arms into resonance.  You can see in the above time series that that's not true.  The left side of the plot is PRMI locked, arms held out of resonance using ALS.  First the Yarm is brought close to resonance, then the Xarm follows.  The RIN of the arms is maybe increasing a little bit as we get closer to resonance, but not by that much.  But there seems to be no correlation between arm power and RIN of the power recycling cavity.

Alternatively, here is some time series when the arm powers got pretty high:

PRMI_2arms_29Oct2013_POPrin_highArmPowers.png


Possible Saturation of Signals:

One possibility for our locklosses of PRMI is that some signal somewhere is saturating, so here are some plots showing that that's not true for the error and control signals for the PRMI:

PRMI_2arms_29Oct2013_LSCcontrolSignals.png

Here, for the exact same time, is a set of time series for every optic except the SRM.  We can see that none of the signals are saturating, and I don't see any big differences for the ITMs or ETMs in the times that the PRMI is locked with high arm powers (center of the x-axis on the plot) and times that the PRMI is not locked, so we don't have high arm powers (edges of the plot - first half second, and last full second).  You can definitely see that the PRM moves much more when the PRMI is locked though, in both pitch and yaw. 

PRMI_2arms_29Oct2013_OpLevs_highArmPowers.png

DCPD signals at the same time:

PRMI_2arms_29Oct2013_DCPDs.png

NB:  These latest 3 plots were created with the getdata script, with arguments "-s 1067163405 -d 7".  It may be a good idea to take some spectra starting at, say 1067163406, 1 second in, and going for ~2 seconds. (It turns out that this is kind of a pain, and I can't convince DTT to give me a sensible spectrum of very short duration....we'll just need to do this live next time around).


Things to think about and investigate:

Why are we losing lock? 

On paper, is the (will the) optical spring a problem once we get high resonance in the arms? 

Spectra of oplevs when we're resonating high arm power.

What is the coupling between 110MHz and 165MHz on the REFL165 PD?  Do we need a stronger bandpass? 

Why are things so shakey when the arm power builds up?

Why do PRCL and MICH have different UGFs when the arms are controlled by ALS vs. ETMs misaligned?

Does QPD for arm transmissions switching work?  Can we then start using TRX and TRY for control?

What is the meaning of the similar features in both transmission signals, and the power recycling cavity?  Power fluctuation in the PRC due to PRM motion? 

ELOG V3.1.3-